# DomainForensics: Exposing Face Forgery across Domains via Bi-directional Adaptation

### v1

@@ I Introduction
During the past few years, large number of DeepFake detection methods [8, 9, 10, 11, 12] have emerged. Trained on the recently proposed large DeepFake datasets, such as FaceForensics++ (FF++) [13] and Celeb-DF [14], these detection methods have shown promising performance. However, these methods fall into the category that the training and testing sets are from the same distribution, e.g., the same type of forgery or the same dataset, which unfortunately limits their practical applications, as there are always new types of forgeries emerging continuously and widespreading to everywhere on various social platforms. These new types of forgeries are very unlikely to have been included in the existing datasets, and thus they are unseen to these detectors, causing significant performance degradation (see Fig. 1 top part). This circumstance gives rise to a big challenge to DeepFake detectors, that is, how to detect constantly emerging new forgeries.
@@
    We propose a new DeepFake detection solution called DomainForensics to handle continuously emerged new forgeries. Different from recent efforts, our method focuses on pushing the detectors to learn the common forgery features across domains, that is, to transfer the forgery knowledge from known forgeries to unseen forgeries, instead of seeking the empirical-common traces on the data level.
@@ II Related Work
In this section, we first present an overview of the existing deepfake detection approaches. We then provide a brief review of upsupervised domain adaptation and discuss the differences between the previous works and our approach.
@@ II-A Deepfake Detection
With the advent of large-scale DeepFake datasets, e.g., [14, 13], DeepFake detection has made significant progress in recent years, e.g., [8, 9, 10, 11, 12, 16, 18, 33, 34, 35, 36, 37]. One challenging problem in this task is how to detect constantly emerging new forgeries. The methods [15, 16, 17, 18, 35, 38] enhance generalization ability by exploring elaborate augmentations on pristine videos, with the aim of covering most of the known forgery types. The limitation of these methods is that the augmentation diversity is restricted to known forgeries. Hence, these methods can hardly handle unknown forgeries. Another vein of methods [19, 20, 21, 10, 22, 39, 40] utilize frequency features to improve generalization ability. However, frequency features can easily be disrupted by post-processing such as compression [41]. Inspired by transfer learning, the methods [23, 24, 25, 37] employ zero-shot and few-shot learning to detect new forgeries. Since zero-shot learning cannot access the samples of new forgeries, its performance gain is severely limited. The few-shot learning needs a small portion of samples and corresponding labels of new forgeries. However, although the video-level label is easily obtained, the face-level label is extremely difficult to obtain in practice.
@@ III-D Bi-directional Adaptation
As aforementioned, to learn domain-common forgery features, we consider building up a solution from the perspective of UDA. Existing general domain adaptation usually considers one-direction to transfer knowledge from the source domain to the target domain. However, such a one-direction adaptation is insufficient in learning transferable knowledge, as it neglects to learn from unlabeled target data, as demonstrated in the previous subsection. Thus, we propose a new bi-directional adaptation strategy, which consists of both forward adaptation and backward adaptation. The forward adaptation stage aims to transfer the knowledge from the source domain to the target domain, just as the existing solutions [45, 48, 49]. However, due to the limitation of such a one-direction adaptation, a portion of target domain knowledge is lost in the transfer, thus hindering the performance of the target domain. To eliminate or mitigate this deficiency, we develop a backward adaptation stage to fine-tune the DeepFake detector on the target domain, while retaining the learned knowledge in the forward adaptation stage. Fig. 3 illustrates the proposed bi-directional adaptation strategy.
@@ III-D2 Backward Adaptation
where $k$ is the class index and $\tau$ is a scalar temperature parameter. $\mathcal{P}^{\prime}(k|x_{i};\tau)$ is obtained by replacing $\mathcal{F}$ with $\mathcal{F}^{\prime}$ in (5). Minimizing this loss helps the student model to excavate more knowledge from the target domain. To achieve the knowledge transfer, we utilize the adversarial loss $\mathcal{L}_{adv}$ as described in the forward adaptation stage on the same discriminator $\mathcal{Q}$ and the distilled feature extractor $\mathcal{F}^{\prime}$. The overall loss of this stage is the combination of these two loss terms as
@@ III-E Network Framework
Fig. 4 depicts the network architecture for our DeepFake detector. We design a ViT-based network as our feature extractor $\mathcal{F}$ due to its strong power on vision tasks. Specifically, our feature extractor has two branches, a visual branch and a frequency branch. For the visual branch, we split the face image into $196$ patches. These patches are flattened and linear transformed to patch embeddings, which are then equipped with position embeddings as the tokens for a ViT [29] architecture to extract visual features. The ViT in this branch contains $l$ transformer layers, each of which is composed by a multi-headed self-attention (MSA) layer and a MLP layer [57]. For the frequency branch, we first convert the face image from RGB color space to YCbCr color space and then apply DCT transformation to each component with a $8\times 8$ block [10]. The transformed frequency maps are concatenated together and sent into a convolution block. We then flatten the feature maps from the convolution block into $196$ vectors, which are used as the input to another ViT architecture for frequency feature extraction. This ViT architecture contains $m$ transformer layers. The visual features and frequency features are concatenated together as the forgery features for DeepFake detection. The classifier $\mathcal{G}$ has a simple structure with only two linear layers, which takes the forgery features as input and outputs the logit of prediction. It is worth emphasizing that our method is independent of the network architecture, and can be integrated into other mainstream architectures. We use the architecture of Fig. 4 in our experiments as it can achieve the best performance, which will be confirmed by Table LABEL:tab:more_archs in Ablation Study.
@@ IV-A1 Datasets
We evaluate our approach on three public deepfake detection datasets, which are FF++ [13], Celeb-DF [14], and StyleGAN [58].
@@
Celeb-DF was proposed more recently, which provides a diverse set of challenges, such as pose, lighting conditions, facial expressions and camera quality, commonly encountered in real-world scenarios. It also offers different levels of manipulation difficulty, ranging from subtle and realistic manipulations to more obvious and noticeable ones. This dataset contains $590$ original videos and 5639 DeepFake videos. We respectively use $5710$ and $518$ videos for training and testing. To construct the training data, we randomly extract $32$ frames from real videos and $4$ frames from fake videos for data balance. As for the testing set, we extract $16$ frames from real and fake videos. The face is cropped out using the same way as for FF++.
StyleGAN [58] trains a generative adversairal network on Flickr-Faces-HQ dataset (FFHQ), which consists of 70000 real face images, to synthesize high-quality fake faces. The synthesizing process focuses on human faces and offers various high-quality synthetic faces by disentangling style and content information in latent space to control the process of face generation. We randomly select 2500 original images from FFHQ and 2500 synthetic images from the generated dataset, where 2000 original images and 2000 synthetic images are for training and the remaining images are for testing.
@@ IV-A2 Domain Adaptation Scenarios
TABLE I: The performance of our proposed method under two cross-manipulations scenarios on FF++. The top part is the results of O2O and the bottom part is the results of O2M.
@@
(HQ) Ours (O2O) 99.08 95.03 99.37 99.77 95.85 99.89 97.44 99.28 97.81 99.14 96.27 98.90
@@
(LQ) Ours (O2O) 92.13 86.89 98.09 93.52 79.64 94.22 86.96 96.38 93.43 88.16 80.85 99.02
(a) Without ours FS $\rightarrow$ F2F
(b) With ours FS $\rightarrow$ F2F
(c) Without ours FS $\rightarrow$ NT
(d) With ours FS $\rightarrow$ NT
@@ IV-A3 Implementation Details
Following the design in Vit-Base [29], the feature extractor contains 12 (i.e., $l=12$) transformer blocks, while we use 4 (i.e., $m=12$) transformer blocks in frequency branch. The input images are resized to $(224,224)$. The weights of loss terms are set to $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=1$. The training batch size is set to $32$ in forward adaptation and $24$ in backward adaptation, as the teacher model occupy extra GPU memory. In the forward adaptation stage, we employ SGD as the optimizer, where the learning rate, momentum, and weight decay are set to $0.001$, $0.9$, and $0.0005$, respectively. We stop the training of forward adaptation after $20$ epochs. For backward adaptation, we set the learning rate to $0.0001$ and the number of training epochs to $10$. All experiments are conducted using Pytorch [60] on two Nvidia GTX 2080Ti GPUs.
@@ IV-B1 Performance of O2O and O2M
In O2O, we select one of the four manipulation methods (DF, F2F, NT, FS) as the source domain and select a different method as the target domain. The top part of Table I shows the performance of our method under this O2O setting on the HQ set. Gray and yellow coloured values denote the performance on the source domain and target domain, respectively. The results reveal that our method performs favourably on both source and target domains, achieving the AUC metric $95\%+$ on most of the source and target adaptation pairs. This demonstrates the effectiveness of our method. We also observe that the performance with NT $\rightarrow$ FS is limited on the source. This is probably due to the large gap between NT and FS, and thus part of the knowledge in the source domain is lost.
@@ IV-B2 Comparison with State-of-the-Arts
We compare our method with several state-of-the-art methods, including the augmentation-based methods (Face X-ray [16], SBI [18]), frequency-based methods (SRM [21], FDFL [10]), transfer-based methods (FOT [23], DDT [24], FEAT [54]), and other methods (Xception [13], MATT [61], LTW [34], RECCE [62], SOLA-sup [63], FTCN [64]).
@@
For the HQ level, it can be seen from the top part of Table II that our method outperforms the augmentation-based methods, Face X-ray and SBI, by a large margin. Since these two methods require prior knowledge of manipulation, they can hardly handle the forgery that has notable differences from the prior knowledge. The frequency-based method SRM performs well on several cases, notably, FS $\rightarrow$ F2F and NT $\rightarrow$ F2F. But its performances are highly degraded in many more cases compared to our method. For example, in DF $\rightarrow$ FS, our method is nearly 100% better than SRM, and in FS $\rightarrow$ DF, our method is 44% better than SRM, while in DF $\rightarrow$ F2F, our method is 30% better than SRM, and in F2F $\rightarrow$ DF, our method is 20% better than SRM. Similar observations can be drawn by comparing our method with SOLA-sup. Based on the available data from [64], the performance of our method are slightly better than those of FTCN. It can also be seen that our method significantly outperforms LTW, Xception, FOT, DDT and FEAT. Moreover, in the more challenging LQ level, our method significantly outperforms all the five benchmarks, as can be seen from the bottom part of Table II.
@@
(LQ) Ours (O2O) 92.13 86.89 98.09 93.52 79.64 94.22 86.96 96.38 93.43 88.16 80.85 99.02
Fig. 5 shows the t-SNE [65] visualizations of feature distribution for two O2O cross-manipulations without and with adaptation using our method. It can be seen that the features of real and fake faces separate well using our method, in contrast to the mixed features without our method. Fig. 6 shows several examples of Grad-CAM [56] (the left six columns), which reveals that our method catches more discriminative features on face regions for both real and fake faces.
@@ IV-B3 Comparison with Existing UDA Methods
Method Training Set Testing Set
@@
(HQ) LTW [34] FF++ 64.10
(HQ) DSP-FWA [15] FF++ 69.30
(HQ) MTD-Net [36] FF++ 70.12
(HQ) Xception [13] FF++ 73.04
(HQ) CFFs [37] FF++ 74.20
(HQ) DAM [67] FF++ 75.30
(HQ) SOLA-sup [63] FF++(DF) 76.02
(HQ) SPSL [68] FF++ 76.88
(HQ) LiSiam [38] FF++ 78.21
(HQ) SLADD [35] FF++ 79.70
(HQ) LipForensics [69] FF++ 82.40
(HQ) HFI-Net [39] FF++ 83.28
(HQ) FTCN [64] FF++ 86.90
[…]
(RAW) PCL+I2G [17] FF++ 90.03
(HQ) Ours FF++ 90.22
[…]
(LQ) MATT [61] FF++ 67.02
(LQ) RECCE [62] FF++ 68.71
(LQ) Two-branch [66] FF++ 73.41
(LQ) Ours FF++ 81.39

### v2

@@ I Introduction
During the past few years, large number of DeepFake detection methods [8, 9, 73, 72, 10, 11, 12] have emerged. Trained on the recently proposed large DeepFake datasets, such as FaceForensics++ (FF++) [13] and Celeb-DF [14], these detection methods have shown promising performance. However, these methods fall into the category that the training and testing sets are from the same distribution, e.g., the same type of forgery or the same dataset, which unfortunately limits their practical applications, as there are always new types of forgeries emerging continuously and widespreading to everywhere on various social platforms. These new types of forgeries are very unlikely to have been included in the existing datasets, and thus they are unseen to these detectors, causing significant performance degradation (see Fig. 1 top part). This circumstance gives rise to a big challenge to DeepFake detectors, that is, how to detect constantly emerging new forgeries.
@@
    We propose a new DeepFake detection solution called DomainForensics to handle continuously emerged new forgeries. Different from recent efforts, our method focuses on pushing the detectors to learn the common forgery features across domains, that is, to transfer the forgery knowledge from known forgeries to unseen forgeries, instead of empirically blending faces on the data level.
@@ II Related Work
In this section, we first present an overview of the existing deepfake detection approaches. We then provide a brief review of unsupervised domain adaptation and discuss the differences between the previous works and our approach.
@@ II-A Deepfake Detection
With the advent of large-scale DeepFake datasets, e.g., [14, 13], DeepFake detection has made significant progress in recent years, e.g., [8, 9, 10, 11, 12, 16, 18, 73, 72, 33, 34, 35, 36, 37]. One challenging problem in this task is how to detect constantly emerging new forgeries. The methods [15, 16, 17, 18, 35, 38, 74] enhance generalization ability by exploring elaborate augmentations on pristine videos, with the aim of covering most of the known forgery types. The limitation of these methods is that the augmentation diversity is restricted to known forgeries. Hence, these methods can hardly handle unknown forgeries. Another vein of methods [19, 20, 21, 10, 22, 39, 40] utilize frequency features to improve generalization ability. However, frequency features can easily be disrupted by post-processing such as compression [41]. Inspired by transfer learning, the methods [23, 24, 25, 37] employ zero-shot and few-shot learning to detect new forgeries. Since zero-shot learning cannot access the samples of new forgeries, its performance gain is severely limited. The few-shot learning needs a small portion of samples and corresponding labels of new forgeries. However, although the video-level label is easily obtained, the face-level label is extremely difficult to obtain in practice.
@@ III-D Bi-directional Adaptation
As aforementioned, to learn domain-common forgery features, we consider building up a solution from the perspective of UDA. Existing general domain adaptation usually considers one-direction to transfer knowledge from the source domain to the target domain. However, such an one-direction adaptation is insufficient in learning transferable knowledge, as it neglects to learn from unlabeled target data, as demonstrated in the previous subsection. Thus, we propose a new bi-directional adaptation strategy, which consists of both forward adaptation and backward adaptation. The forward adaptation stage aims to transfer the knowledge from the source domain to the target domain, just as the existing solutions [45, 48, 49]. However, due to the limitation of such an one-direction adaptation, a portion of target domain knowledge is lost in the transfer, thus hindering the performance of the target domain. To eliminate or mitigate this deficiency, we develop a backward adaptation stage to fine-tune the DeepFake detector on the target domain, while retaining the learned knowledge in the forward adaptation stage. Fig. 3 illustrates the proposed bi-directional adaptation strategy.
@@ III-D2 Backward Adaptation
where $k$ is the class index and $\tau$ is a scalar temperature parameter. $\mathcal{P}^{\prime}(k|x_{i};\tau)$ is obtained by replacing $\mathcal{F}$ with $\mathcal{F}^{\prime}$ in (5). The Eq.(4) is designed to distill the knowledge from the teacher model learned in the target domain into the student model. By using self-distillation, the student model can enhance the features of the teacher model, thus capturing more effective knowledge regarding the target domain in comparison to the teacher model. We then utilize $\mathcal{L}_{adv}$ as described in the forward adaptation stage on the same discriminator $\mathcal{Q}$ and the distilled feature extractor $\mathcal{F}^{\prime}$. The overall loss of this stage is the combination of these two loss terms as
@@ III-E Network Framework
Fig. 4 depicts the network architecture for our DeepFake detector. We design a ViT-based network as our feature extractor $\mathcal{F}$ due to its strong power on vision tasks. Specifically, our feature extractor has two branches, a visual branch and a frequency branch. For the visual branch, we split the face image into $196$ patches. These patches are flattened and linear transformed to patch embeddings, which are then equipped with position embeddings as the tokens for a ViT [29] architecture to extract visual features. The ViT in this branch contains $l$ transformer layers, each of which is composed by a multi-headed self-attention (MSA) layer and a MLP layer [57]. For the frequency branch, we first convert the face image from RGB color space to YCbCr color space and then apply DCT transformation to each component with a $8\times 8$ block [10]. The transformed frequency maps are concatenated together and sent into a convolution block. We then flatten the feature maps from the convolution block into $196$ vectors, which are used as the input to another ViT architecture for frequency feature extraction. This ViT architecture contains $m$ transformer layers. The visual features and frequency features are concatenated together as the forgery features for DeepFake detection. The classifier $\mathcal{G}$ has a simple structure with only two linear layers, which takes the forgery features as input and outputs the logit of prediction. It is worth emphasizing that our method is independent of the network architecture, and can be integrated into other mainstream architectures. We use the architecture of Fig. 4 in our experiments as it can achieve the best performance, which will be confirmed by Table IX in Ablation Study.
@@ IV-A1 Datasets
We evaluate our approach on three public deepfake detection datasets, which are FF++ [13], Celeb-DF [14], StyleGAN [58], DFDCP, and FFIW.
@@
Celeb-DF was proposed more recently, which provides a diverse set of challenges, such as pose, lighting conditions, facial expressions and camera quality, commonly encountered in real-world scenarios. It also offers different levels of manipulation difficulty, ranging from subtle and realistic manipulations to more obvious and noticeable ones. This dataset contains $590$ original videos and $5,639$ DeepFake videos. We respectively use $5,710$ and $518$ videos for training and testing. To construct the training data, we randomly extract $32$ frames from real videos and $4$ frames from fake videos for data balance. As for the testing set, we extract $16$ frames from real and fake videos. The face is cropped out using the same way as for FF++.
StyleGAN [58] trains a generative adversairal network on Flickr-Faces-HQ dataset (FFHQ), which consists of $70,000$ real face images, to synthesize high-quality fake faces. The synthesizing process focuses on human faces and offers various high-quality synthetic faces by disentangling style and content information in latent space to control the process of face generation. We randomly select $2,500$ original images from FFHQ and $2,500$ synthetic images from the generated dataset, where $2,000$ original images and $2,000$ synthetic images are for training and the remaining images are for testing.
DFDCP [71] is a preview dataset for deepfakes detection challenge consisting of $5,000$ videos with two facial modification algorithms, considering diversity in several axes (gender, skin-tone, age, etc.).
FFIW-10K [67] is a large-scale dataset, which comprises $10,000$ pristine videos and $10,000$ high-quality forgery videos for training. Another $250$ real videos and $250$ manipulated videos are also provided for testing. Different from the aforementioned datasets, this dataset contains multiple forgery faces in a single video.
@@ IV-A2 Domain Adaptation Scenarios
TABLE I: The performance of our proposed method under two cross-manipulation scenarios on FF++. The top part is the results of O2O and the bottom part is the results of O2M.
@@
(HQ) TALL-Swin [73] 52.96 52.11 77.36 62.55 60.49 55.54 64.92 62.09 52.20 50.55 51.96 85.08
(HQ) Baseline 69.43 69.55 55.18 81.71 55.84 69.74 55.19 79.74 44.98 71.06 43.63 84.34
(HQ) Baseline+BA (Ours) 99.08 95.03 99.37 99.77 95.85 99.89 97.44 99.28 97.81 99.14 96.27 98.90
@@
(LQ) TALL-Swin [73] 55.50 52.93 74.39 63.64 57.82 53.15 68.65 63.19 51.42 50.44 53.11 83.91
(LQ) Baseline 61.18 59.49 71.17 65.88 51.43 59.07 56.36 65.33 52.49 56.25 44.38 70.45
(LQ) Baseline+BA (Ours) 92.13 86.89 98.09 93.52 79.64 94.22 86.96 96.38 93.43 88.16 80.85 99.02
(a) Without ours FS $\rightarrow$ F2F (F2F)
(b) With ours FS $\rightarrow$ F2F (F2F)
(c) Without ours FS $\rightarrow$ NT (NT)
(d) With ours FS $\rightarrow$ NT (NT)
(e) Without ours FS $\rightarrow$ F2F (FS)
(f) With ours FS $\rightarrow$ F2F (FS)
(g) Without ours FS $\rightarrow$ NT (FS)
(h) With ours FS $\rightarrow$ NT (FS)
@@ IV-A3 Implementation Details
Following the design in Vit-Base [29], the feature extractor contains 12 (i.e., $l=12$) transformer blocks, while we use 4 (i.e., $m=12$) transformer blocks in frequency branch. The input images are resized to $(224,224)$. The weights of loss terms are set to $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=1$. The training batch size is set to $32$ in forward adaptation and $24$ in backward adaptation, as the teacher model occupy extra GPU memory. In the forward adaptation stage, we employ SGD as the optimizer, where the learning rate, momentum, and weight decay are set to $0.001$, $0.9$, and $0.0005$, respectively. We stop the training of forward adaptation after $20$ epochs. For backward adaptation, we set the learning rate to $0.0001$ and the number of training epochs to $10$. The temperature is set to $\tau=0.5$. All experiments are conducted using Pytorch [60] on two Nvidia GTX 2080Ti GPUs.
@@ IV-B1 Performance of O2O and O2M
In O2O, we select one of the four manipulation methods (DF, F2F, NT, FS) as the source domain and select a different method as the target domain. The top part of Table I shows the performance of our method under this O2O setting on the HQ set. Gray and yellow coloured values denote the performance on the source domain and target domain, respectively. “Baseline” denotes training the proposed architecture on source domain and directly testing on other domains without bi-directional adaptation. The results reveal that our method performs favourably on both source and target domains, achieving the AUC metric $95\%+$ on most of the source and target adaptation pairs. This demonstrates the effectiveness of our method. We also observe that the performance with NT $\rightarrow$ FS is limited on the source. This is probably due to the large gap between NT and FS, and thus part of the knowledge in the source domain is lost.
@@ IV-B2 Comparison with State-of-the-Arts
We compare our method with several state-of-the-art methods, including the augmentation-based methods (Face X-ray [16], SBI [18]), frequency-based methods (SRM [21], FDFL [10]), transfer-based methods (FOT [23], DDT [24], FEAT [54]), and other methods (Xception [13], MATT [61], LTW [34], RECCE [62], SOLA-sup [63], FTCN [64], TALL-Swin [73]). Note that TALL-Swin has no reports corresponding to this scenario. Thus we reproduce its results following its default settings. This method is trained using all videos in FF++.
@@
For the HQ level, it can be seen from the top part of Table II that our method outperforms the augmentation-based methods, Face X-ray and SBI, by a large margin. Since these two methods require prior knowledge of manipulation, they can hardly handle the forgery that has notable differences from the prior knowledge. The frequency-based method SRM performs well on several cases, notably, FS $\rightarrow$ F2F and NT $\rightarrow$ F2F. But its performances are highly degraded in many more cases compared to our method. For example, in DF $\rightarrow$ FS, our method is nearly 100% better than SRM, and in FS $\rightarrow$ DF, our method is 44% better than SRM, while in DF $\rightarrow$ F2F, our method is 30% better than SRM, and in F2F $\rightarrow$ DF, our method is 20% better than SRM. Similar observations can be drawn by comparing our method with SOLA-sup. Based on the available data from [64], the performance of our method are slightly better than those of FTCN. When compared to TALL-Swin, we observe our method achieves better generalization performances than TALL-Swin. It can also be seen that our method significantly outperforms LTW, Xception, FOT, DDT and FEAT. Moreover, in the more challenging LQ level, our method significantly outperforms all the five benchmarks, as can be seen from the bottom part of Table II.
@@
(LQ) Ours 92.13 86.89 98.09 93.52 79.64 94.22 86.96 96.38 93.43 88.16 80.85 99.02
Fig. 5 shows the t-SNE [65] visualizations of feature distribution for two O2O cross-manipulations (FS $\rightarrow$ F2F, FS $\rightarrow$ NT) without and with adaptation using our method on both source and target domain. For the target domain, it can be seen that the features of real and fake faces separate well using our method, in contrast to the mixed features without our method (see (a,b) and (c,d)). For the source domain, we can observe that their feature distributions are still discriminative (see (e,f) and (g,h)). Fig. 6 shows several examples of Grad-CAM [56] (the left six columns), which reveals that our method catches more discriminative features on face regions for both real and fake faces.
@@ IV-B3 Comparison with Existing UDA Methods
Method Training Set FF++ Testing Set
@@
(HQ) LTW [34] FF++ - 64.10
(LQ) MATT [61] FF++ 99.80 67.02
(HQ) DSP-FWA [15] FF++ - 69.30
(HQ) MTD-Net [36] FF++ 99.38 70.12
(HQ) Xception [13] FF++ 95.73 73.04
(HQ) CFFs [37] FF++ 97.03 74.20
(HQ) DAM [67] FF++ - 75.30
(HQ) SOLA-sup [63] FF++(DF) - 76.02
(HQ) SPSL [20] FF++ 95.32 76.88
(HQ) LiSiam [38] FF++ 99.13 78.21
(HQ) SLADD [35] FF++ 98.40 79.70
(HQ) LipForensics [68] FF++ 99.70 82.40
(HQ) HFI-Net [39] FF++ 98.86 83.28
(HQ) TALL-ViT [73] FF++ - 86.58
(HQ) TALL-Swin [73] FF++ - 90.79
(HQ) FTCN [64] FF++ - 86.90
(HQ) AltFreezing [72] FF++ 99.70 89.50
(HQ) F2Trans [40] FF++ 99.74 89.87
(RAW) PCL+I2G [17] FF++ 99.79 90.03
(HQ) Baseline FF++ 99.13 81.90
(HQ) Baseline+BA (Ours) FF++ 98.64 90.22
(LQ) F3-Net [19] FF++ 93.30 61.51
(LQ) RECCE [62] FF++ 95.02 68.71
(LQ) Two-branch [66] FF++ 93.18 73.41
(LQ) Baseline FF++ 94.25 75.68
(LQ) Baseline+BA (Ours) FF++ 91.24 81.39